OpenStudy (anonymous):

What is the surface area of this rectangular prism? Rectangular prism with length labeled 11 feet, height labeled 2 feet, and width labeled 6 feet. ft²

OpenStudy (anonymous):



OpenStudy (anonymous):

@unkownbeast911 can u help

OpenStudy (anonymous):

SA= 2(lw)+2(lh)+2(hw) SA= 2(22)+2(12)+2(66) SA= 44+24+132 SA= 200 ft²

OpenStudy (anonymous):

wow so wats the answer

OpenStudy (anonymous):

200 ft²

OpenStudy (anonymous):

thxs here yo medal

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

can u help with 1 more

OpenStudy (anonymous):

Look at the cube, and the net of this cube. What is the surface area of this cube?

OpenStudy (anonymous):



OpenStudy (anonymous):

@michelle87 help

OpenStudy (anonymous):

864 mm²

OpenStudy (anonymous):

ok show me how u got it plz

OpenStudy (anonymous):

SA= 2(lw)+2(lh)+2(hw) SA= 2(144)+2(144)+2(144) SA= 288+288+288 SA= 864 mm²

OpenStudy (anonymous):

wow ok